A 48% glyphosate solution contains 2.84 mol of glyphosate per liter. What is the molecular weight of glyhosate?
Answer #1

Let M be the molecular weight of glyphosate.

Number of moles = given weight/ molecular weight

=> Given weight = Number of moles * M = 2.84 * M

It is given that 48% glyphosate solution caontains 2.84 mol per liter of glyphosate.

=> 0.48 = weight of glyphosate/weight of water

=> 0.48 = (2.84 * M)/1000 (2.84 is already in per liter units and 1 liter of water ~ 1000 grams)

=> M = (0.48*1000)/2.84 = 480/2.84 = 169.01 grams

Hence the glyphosate has molecular weight of 169 gram/mol

Hope this helped!